# Vulnerability Coordination Under the Cyber Resilience Act

### v1

@@ 1 Introduction
The CRA is a complex piece of legislation covering practically the whole information technology sector. Both hardware and software products are in its scope. Among the notable exemptions not covered by the regulation, as specified in the CRA’s Article 2, are medical devices, motor vehicles, and ships and marine equipment. Furthermore, recital 12 clarifies that cloud computing is also excluded, including software as a service (SaaS), platform as a service (PaaS), and infrastructure as a service (IaaS) solutions to which the NIS2 directive applies instead. In general, the CRA categorizes products into two categories: low-risk and high-risk products. Much more requirements are imposed upon the high-risk products, including mandatory security audits conducted by third-parties. However, in what follows, the focus is only on the CRA’s implications for vulnerability coordination, including vulnerability disclosure and associated reporting. With this focus in mind, the opening Section 2 introduces the vulnerability coordination and disclosure concepts. Then, the subsequent Section 3 discusses and elaborates the CRA’s new implications to vulnerability coordination and disclosure. The implications are not uniform across everyone; therefore, Section 4 continues by briefly elaborating the requirements for open source software products and their projects. To some extent, the CRA addresses also the pressing issues with supply chain security [31]. The important supply chain topic is discussed in Section 5. A conclusion and a discussion are presented in final Section 6.
@@ 2 Vulnerability Disclosure and Coordination
For this reason and other associated reasons, a third vulnerability disclosure model emerged. It is called coordinated vulnerability disclose; in essence, a discoverer contacts an intermediary party who helps with the coordination and collaboration with a vendor affected, often also providing advices to other parties and the general public through security advisories and security awareness campaigns. This coordinated disclosure model is the de facto vulnerability disclosure practice today. Regarding the intermediaries involved, a national CSIRT is typically acting as a coordinator, but also commercial companies may act in a similar role, as is typical in the nowadays popular so-called bug bounties and their online platforms. In addition to security improvements, commercial bug bounty platforms may give vendors more control over disclosure processes [1], while for discoverers they offer monetary rewards, learning opportunities, and some legal safeguards [2]. It is also important to emphasize that today practically all coordinated vulnerability disclosure models further follow a so-called responsible disclosure, meaning that vendors are given a fixed time period to develop and release patches before information is released to the public.
@@ 3 The CRA’s New Requirements
Regarding vulnerabilities, as can be seen also from Fig. 1, it should be underlined that the mandatory notifications only involve actively exploited vulnerabilities. Thus, among other things, so-called silent patching is excluded from the CRA’s scope; it refers to cases in which a vendor has found a vulnerability from its product and silently corrected it without making the information available to others [26]. A more important point is that the definition for actively exploited vulnerabilities is rather loose. When considering a real possibility of underreporting—a topic that has been discussed in the literature in somewhat similar settings [14], particularly the wording about reliable evidence raises a concern. A relevant question is: whose evidence? If vendors alone answer to the question, underreporting may well occur. Though, other parties too, including public authorities, have threat intelligence platforms, service integration and management (SIEM) systems, and many related techniques deployed. Against this backdrop, a vendor deciding not to report an actively exploited vulnerability may later on find itself as having been non-compliant. In addition, Article 60 allows market surveillance authorities to conduct sweeps to check compliance.
These points are important because the CRA implies potentially large administrative fines from non-compliance; Article 64 specifies the maximum administrative fine to be either ten million euros or 2% of the total worldwide annual turnover of a vendor, whichever is higher. Furthermore, Aticle 58(2) specifies that the member states may both prohibit a severely non-compliant product from entering the EU’s whole internal market and withdraw severely non-complaint products already sold on the internal market. These prohibition and withdrawal rights resemble those present in the EU’s older product safety laws [24]. Analogously to product safety, it may be that these rights granted to market surveillance authorities will act as an additional deterrent against non-compliance.
@@ 6 Conclusion
Of the new regulatory obligations and points raised thereto, (1) further research is required about the concept of actively exploited vulnerabilities. Such research might also generally reveal whether the existing empirical vulnerability research has been biased in a sense that serious vulnerabilities affecting actual deployments have perhaps been downplayed; with some notable exceptions [4], the focus has often been on reported vulnerabilities affecting software products, components, and vendors, not on their actual deployments [20]. On a more theoretical side of research, (2) further contributions are required also regarding the other fundamental concepts involved, including not only vulnerabilities but also security incidents and the so-called near misses. Then, regarding supply chains, (3) further research is required also on SBOMs. Although technical research is well underway in this regard, including with respect to a question of how to generate SBOMs automatically [22], many questions remain about applicability, usability, and related aspects, including for regulatory purposes. Although perhaps only partially related to the paper’s explicit theme, (4) a relevant research question would also be whether, how, and how well vendors can comply with the requirement to provide security patches, preferably in an automated manner. When recalling the CRA’s scope, the question is non-trivial. While providing automated security updates for the so-called Internet of things, or the “Internet of trash”, as it has sometimes been called [17], sounds like a reasonable and well-justified goal, the situation may be quite different in industrial settings. For instance, patching an industrial robot, whether for security or other reasons, may involve also safety considerations, while more generally updating a kernel, a firmware, or any component thereto may involve serious risks, as CrowdStrike found out in 2024.111 https://en.wikipedia.org/wiki/2024_CrowdStrike-related_IT_outages It also remains generally unclear whether the CRA applies to physical security, including patching of software operating in air-gapped and related deployments. As always, (5) further policy-oriented research is also needed. Among other things, a close eye is needed to monitor and evaluate the regulation’s future implementation, adoption, administration, and enforcement.
@@
Regarding the CRA’s future implementation, (6) a particularly relevant research topic involves examining the upcoming technical standards upon which the essential cyber security requirements are based. This topic is important already because complying with these standards provides a presumption of conformity according to the CRA’s Article 27(1). The topic can be extended to other (7) conformity assessment procedures involved, including the mandatory third-party audits required for high-risk products and the CRA’s relation to the older cyber security certification schemes in the EU. Among other things, a good topic to examine would be the accreditation of conformity assessment bodies. It may be that the CRA will also entail a new labor market demand for cyber security professionals specialized into security testing, auditing, and related domains. At the same time, as also implicitly remarked in the CRA’s Article 10, there is allegedly a cyber security skill gap in Europe and elsewhere [13]. As the gap may correlate with education [23, 32], (8) it may also be worthwhile to further study the CRA and its future implementation with education and curricula in mind. Regarding adoption, (9) it is relevant to know which conformity assessment procedures vendors will prefer in the future; according to Article 32(1), there are four distinct procedures for conformity assessments under the CRA.
@@
Although it is too early to speculate what the CRA’s broader implications may be, four general points can be still raised. First, the larger regulatory cyber security package in the EU has expectedly raised some criticism that fragmentation and complexity have increased [9, 25]. The CRA has contributed to this increase. The CRA also interacts with the EU’s other (cyber security) laws. In particular, as the NIS2 directive also imposes reporting obligations for vendors, synchronization between the CRA and NIS2 has been perceived as crucial [7]. Second, a critical reflection has been presented regarding the potential offered by the so-called Brussels effect on one hand and the potential negative economic consequences on the other [31]. However, it should be emphasized that there are numerous other factors behind this trade-off conundrum; the CRA is only a small piece of a large puzzle in this regard. Although mileages have varied [30], and the EU has established funding instruments for implementing its new cyber security laws [25], it nevertheless remains unclear whether sufficient investments have been made, as it remained ten years ago [21]. Third, a notable limitation of the CRA is that it offers no legal guards against criminal or civil liability from disclosing vulnerabilities. Such guards have often been seen important in the literature [2, 5]. Last but not least, it remains to be seen whether the CRA actually improves the cyber security of products. Although all points require further research, including rigorous evaluations, particularly the last point is fundamental in terms of the CRA’s explicit goals. According to the Act’s first recital, cyber insecurity affects not only the EU’s economy but also its democracy as well as the safety and health of consumers. Already against this backdrop, it is important and relevant to evaluate in the future how well the CRA managed to remove or at least mitigate the overall cyber insecurity in the European Union.

### v2

@@ 1 Introduction
The CRA covers the whole information technology sector and beyond. Both hardware and software products with a networking functionality are in its scope. Among the notable exemptions not covered by the regulation, as specified in the CRA’s Article 2, are medical devices, motor vehicles, and ships and marine equipment. Furthermore, recital 12 clarifies that cloud computing is also excluded. The CRA categorizes products into three categories: “normal”, “important”, and “critical” products. While all products need to comply with the CRA’s essential requirements, more obligations are placed upon the important and critical products [24]. However, in what follows, the focus is only on the CRA’s implications for vulnerability coordination, including vulnerability disclosure. With this focus in mind, the opening Section 2 introduces the vulnerability coordination and disclosure concepts. Then, the subsequent Section 3 discusses and elaborates the CRA’s new implications to vulnerability coordination and disclosure. The implications are not uniform across everyone; therefore, Section 4 continues by briefly elaborating the requirements for open source software products and their projects. To some extent, the CRA addresses also the pressing issues with supply chain security. The important supply chain topic is discussed in Section 5. A conclusion and a discussion are presented in the final Section 6.
@@ 2 Vulnerability Disclosure and Coordination
For this reason and other associated reasons, a third vulnerability disclosure model emerged. It is called coordinated vulnerability disclosure; in essence, a discoverer contacts an intermediary party who helps with the coordination and collaboration with a vendor affected, often also providing advices to other parties and the general public through security advisories and security awareness campaigns. This coordinated disclosure model is the de facto vulnerability disclosure practice today. Regarding the intermediaries involved, a national CSIRT is typically acting as a coordinator, but also commercial companies may act in a similar role, as is typical in the nowadays popular so-called bug bounties and their online platforms. In addition to security improvements, commercial bug bounty platforms may give vendors more control over disclosure processes [1], while for discoverers they offer monetary rewards, learning opportunities, and some legal safeguards [2]. It is also important to emphasize that today practically all coordinated vulnerability disclosure models further follow a so-called responsible disclosure, meaning that vendors are given a fixed time period to develop and release patches before information is released to the public.
@@ 3 The CRA’s New Requirements
Regarding vulnerabilities, as can be seen also from Fig. 1, it should be underlined that the mandatory notifications only involve actively exploited vulnerabilities. Thus, among other things, so-called silent patching is excluded from the CRA’s scope; it refers to cases in which a vendor has found a vulnerability from its product and silently corrected it without making the information available to others [25]. A more important point is that the definition for actively exploited vulnerabilities is rather loose. When considering a possibility of underreporting—a topic that has been discussed in the literature in somewhat similar settings [14], particularly the wording about reliable evidence raises a concern. A relevant question is: whose evidence? If vendors alone answer to the question, underreporting may occur. Though, other parties too, including public authorities, have threat intelligence platforms, security information and event management (SIEM) systems, and many related techniques deployed. Against this backdrop, a vendor deciding not to report an actively exploited vulnerability may later on find itself as having been non-compliant. In addition, Article 60 allows market surveillance authorities to conduct sweeps to check compliance.
@@ 6 Conclusion
Of the new regulatory obligations and points raised thereto, (1) further research is required about the concept of actively exploited vulnerabilities. Such research might also generally reveal whether the existing empirical vulnerability research has been biased in a sense that serious vulnerabilities affecting actual deployments have perhaps been downplayed; with some notable exceptions [4], the focus has often been on reported vulnerabilities affecting software products, components, and vendors, not on their actual deployments [19]. On a more theoretical side of research, (2) further contributions are required also regarding the other fundamental concepts involved, including not only vulnerabilities but also security incidents and the so-called near misses. Then, regarding supply chains, (3) further research is required also on SBOMs. Although technical research is well underway in this regard, including with respect to a question of how to generate SBOMs automatically [20], many questions remain about applicability, usability, and related aspects, including for regulatory purposes. Although perhaps only partially related to the paper’s explicit theme, (4) a relevant research question would also be whether, how, and how well vendors can comply with the requirement to provide security patches, preferably in an automated manner. When recalling the CRA’s scope, the question is non-trivial. While providing automated security updates for Internet of things devices sounds like a reasonable and well-justified goal, the situation may be quite different in industrial settings [24]. It also remains generally unclear whether the CRA applies to physical security, including patching of software operating in air-gapped and related deployments. As always, (5) further policy-oriented research is also needed. Among other things, a close eye is needed to monitor and evaluate the regulation’s future implementation, adoption, administration, and enforcement.
@@
Regarding the CRA’s future implementation, (6) a particularly relevant research topic involves examining the upcoming technical standards upon which the essential cyber security requirements are based. This topic is important already because complying with these standards provides a presumption of conformity according to the CRA’s Article 27(1). The topic can be extended to other (7) conformity assessment procedures involved, including the mandatory third-party audits required for critical products and the CRA’s relation to the older cyber security certification schemes in the EU. Among other things, a good topic to examine would be the accreditation of conformity assessment bodies. It may be that the CRA will also entail a new labor market demand for cyber security professionals specialized into security testing, auditing, and related domains. At the same time, as also implicitly remarked in the CRA’s Article 10, there is allegedly a cyber security skill gap in Europe and elsewhere [13]. As the gap may correlate with education [21, 29], (8) it may also be worthwhile to further study the CRA and its future implementation with education and curricula in mind. Regarding adoption, (9) it is relevant to know which conformity assessment procedures vendors will prefer in the future; according to Article 32(1), there are four distinct procedures for conformity assessments under the CRA.
@@
Although it is too early to speculate what the CRA’s broader implications may be, three general points can be still raised. First, the larger regulatory cyber security package in the EU has expectedly raised some criticism that fragmentation and complexity have increased [9, 23]. The CRA has contributed to this increase. The CRA also interacts with the EU’s other(cyber security laws. In particular, as the NIS2 directive also imposes reporting obligations for vendors, synchronization between the CRA and NIS2 has been perceived as crucial [7]. Second, a notable limitation of the CRA is that it offers no legal guards against criminal or civil liability from disclosing vulnerabilities. Such guards have often been seen important in the literature [2, 5]. Third, it remains to be seen whether the CRA actually improves the cyber security of products. Although all points require further research, including rigorous evaluations, particularly the last point is fundamental in terms of the CRA’s explicit goals. According to the Act’s first recital, cyber insecurity affects not only the EU’s economy but also its democracy as well as the safety and health of consumers. Already against this backdrop, it is important and relevant to evaluate in the future how well the CRA managed to remove or at least mitigate cyber insecurity in the EU.